OpenStudy (anonymous):

Profit Margin-A car dealer sells Wilsons, which have a sticker price of $9,779. Her usual profit margin on this model is 12%. She wants to increase her profit by $391.16. What would her new profit margin be? 15.4% 19.6% 16.3% I've gotten to the point where i know $1564.64 is the amount she wants and im pretty sure i need to times her initial cost by each percent to get another number to add on to see if it makes the price she wants, but im not sure how to factor in the percents since they are not whole, am i on the right track?

jimthompson5910 (jim_thompson5910):

you don't need to know the final cost of the product all you do is divide the profit margin by the sticker price then you multiply by 100 to get a percentage

jimthompson5910 (jim_thompson5910):

oh wait, nvm, missed a big step one sec

OpenStudy (anonymous):

Hmm i got 81491 and then i timed by 100 and got 8149166

jimthompson5910 (jim_thompson5910):

her profit margin is 12% of the sticker price so 12% of 9779 = 0.12*9779 = 1,173.48 this is her old profit

jimthompson5910 (jim_thompson5910):

she wants to bump this profit up by 391.16

jimthompson5910 (jim_thompson5910):

1,173.48+391.16 = 1,564.64

jimthompson5910 (jim_thompson5910):

so this is the new profit she wants, 1,564.64 dollars now you divide this by the sticker price and multiply by 100 1,564.64/9779 = 0.16 0.16*100 = 16% profit margin

jimthompson5910 (jim_thompson5910):

hmm 16% isn't on the list...closest thing is 16.3%

OpenStudy (anonymous):

Thank you so much ! Your very good with this stuff, I will try 16.3

jimthompson5910 (jim_thompson5910):

thanks, not good enough lol not sure why 16% isn't on the list, i might be missing something?

jimthompson5910 (jim_thompson5910):

can you double check all the numbers and make sure there are no typos?

OpenStudy (anonymous):

Im not sure either it is online practice for summer school ive been having trouble with, only those 3 15.4% 19.6% 16.3% :/

OpenStudy (anonymous):

3 answers *

jimthompson5910 (jim_thompson5910):

i meant the other numbers (like 9779)

jimthompson5910 (jim_thompson5910):

stuff like that

OpenStudy (anonymous):

Oh! yes i can!

jimthompson5910 (jim_thompson5910):

thanks

OpenStudy (anonymous):

No, thank you!

OpenStudy (anonymous):

I checked it all and got the same, 16%, at first i wasn't sure what you meant thats why i took so long, sorry!

jimthompson5910 (jim_thompson5910):

that's ok, thanks for going back and double checking for me you'll have to bring it up with your teacher and s/he will know how to fix it

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

I will do that